# Process vs. Outcome Reward: Which is Better for Agentic RAG Reinforcement Learning

### v2

@@ Abstract
Retrieval-augmented generation (RAG) enhances the text generation capabilities of large language models (LLMs) by integrating external knowledge and up-to-date information. However, traditional RAG systems are limited by static workflows and lack the adaptability required for multistep reasoning and complex task management. To address these limitations, agentic RAG systems (e.g., DeepResearch) have been proposed, enabling dynamic retrieval strategies, iterative context refinement, and adaptive workflows for handling complex search queries beyond the capabilities of conventional RAG. Recent advances, such as Search-R1, have demonstrated promising gains using outcome-based reinforcement learning, where the correctness of the final answer serves as the reward signal. Nevertheless, such outcome-supervised agentic RAG methods face challenges including low exploration efficiency, gradient conflict, and sparse reward signals. To overcome these challenges, we propose to utilize fine-grained, process-level rewards to improve training stability, reduce computational costs, and enhance efficiency. Specifically, we introduce a novel method ReasonRAG that automatically constructs RAG-ProGuide, a high-quality dataset providing process-level rewards for (i) query generation, (ii) evidence extraction, and (iii) answer generation, thereby enhancing model inherent capabilities via process-supervised reinforcement learning. With the process-level policy optimization, the proposed framework empowers LLMs to autonomously invoke search, generate queries, extract relevant evidence, and produce final answers. Compared to existing approaches such as Search-R1 and traditional RAG systems, ReasonRAG, leveraging RAG-ProGuide, achieves superior performance on five benchmark datasets using only 5k training instances, significantly fewer than the 90k training instances required by Search-R1. Our code is available at https://github.com/wlzhang2020/ReasonRAG.
@@ 1 Introduction
Large language models (LLMs) brown2020language ; chowdhery2023palm ; achiam2023gpt have demonstrated substantial proficiency in text generation and natural language understanding tasks zhularge ; gonzalez2024does . However, their reliance on static training data constrains their ability to address dynamic and real-time queries, often resulting in outdated or hallucinated information ji2023survey ; huang2025survey ; mckenna2023sources . Retrieval-Augmented Generation (RAG) lewis2021RAG has emerged as a promising solution by equipping LLMs with external knowledge sources, improving the relevance, factual accuracy, and timeliness of responses gao2023retrieval ; fan2024survey ; xiong2024benchmarking . Despite these advances, traditional RAG architectures are limited by their linear and static workflows, which suffer from complex multi-step reasoning, deep contextual integration, and iterative response refinement lewis2021RAG ; singh2025agentic . To address these shortcomings, agentic RAG (e.g., DeepResearch ds2024gemini ; ds2025openai ) systems have been developed, enabling dynamic retrieval strategies, enhanced contextual understanding, and iterative refinement. Achieving agentic RAG requires the underlying LLMs to orchestrate retrieval, filter relevant information, and iteratively refine their outputs, resulting in more adaptive and efficient information processing.
@@ 2.2.1 Shortest Path Reward Estimation (SPRE)
Formally, the agentic RAG process consists of an $n$-step sequence $[y_{1},\cdots,y_{n}]$, where each $y_{i}$ represents the output of a single reasoning step, conditioned on the initial question $x$ and previous steps $y_{<i}$. To evaluate a partial sequence $y_{1:t}$, we simulate multiple continuations, known as rollouts, until a final answer is obtained. By repeating the rollout process $k$ times and scoring each outcome, we approximate the reward as a Monte Carlo-style estimation with step-based penalties:
@@ 2.2.2 Monte Carlo Tree Search (MCTS) for Process-level Exploration
2.2.3 RAG-ProGuide Dataset
Construction.
Using the MCTS-based exploration framework, we construct a high-quality process-supervised dataset RAG-ProGuide to facilitate process-level policy optimization. We randomly sample 3,000 questions each from PopQA mallen2022not , HotpotQA yang2018hotpotqa , and 2WikiMultihopQA ho2020constructing , covering both single-hop and multi-hop question answering tasks. GPT-4o hurst2024gpt serves as the policy model to simulate the agentic RAG reasoning process within the MCTS framework (see Section 2.2.2). During tree search, we prune all branches that do not yield a final answer. For each complete trajectory, we compute the F1 score between the predicted answer and ground truth, and use this correctness signal to estimate intermediate node rewards via SPRE (see Section 2.2.1). These rewards are propagated through the MCTS tree to guide preference pair selection. To ensure high-quality preference data, we perform post-processing to remove duplicates and uninformative comparisons: (i) we discard preference pairs with identical response sequences, and (ii) pairs with a reward difference less than 0.01. After filtering, the final dataset consists of 4,603 questions and 13,289 distinct preference pairs.
Dataset Statistics.
Table 1 presents detailed statistics and distributions for our dataset. The questions are drawn from PopQA, HotpotQA, and 2WikiMultihopQA, providing comprehensive coverage of both single-hop and multi-hop reasoning within the RAG decision space. The dataset contains a balanced distribution of three reasoning actions, reflecting the multi-stage nature of the agentic RAG process. As shown in Figure 3(a), the distribution of preference pair types demonstrates diverse comparative scenarios; the x-axis abbreviations (A: answer generation, Q: query generation, E: evidence extraction) indicate action types in accepted versus rejected paths. This diversity ensures fine-grained comparative coverage across different reasoning stages. Figure 3(b) indicates a range of reasoning iteration counts, consistent with the complexity of multi-hop inference. Figure 3(c) shows a broad distribution of response token lengths, confirming the dataset’s capacity to capture various response complexities. Additionally, Figure 3(d) highlights substantial reward differences between preference pairs, which is crucial for effective preference learning. Collectively, these statistics demonstrate the dataset’s quality and its suitability for training robust process-level decision policies in agentic RAG frameworks.
@@ 2.3 Process-Supervised Preference Optimization
[…]
2:while $i<N$ do
@@
19:     $i\leftarrow i+1$
20:end while
21:return final response $y$.
@@ 2.4 Agentic RAG Inference
The workflow operates through three recursive decision states: Reasoning, Grounding, and Terminal. In the Reasoning state, the LLM evaluates the current context to decide if it has sufficient information to answer the question. If sufficient, it generates a final answer enclosed in placeholders (<answer> and </answer>), thus terminating the process. If not, the model creates a new query enclosed in <query> and </query> placeholders to retrieve additional evidence. The system then transitions to the Grounding state, where documents are retrieved based on the query, and the model extracts relevant evidence spans. These evidence spans are appended to the context, after which the process loops back to the Reasoning state for further deliberation. (See Appendix G for detailed prompt designs.)
@@ 3.1 Experimental Setup
Evaluation Dataset & Metrics. We evaluate ReasonRAG and all baselines on five public benchmarks: the single-hop QA dataset PopQA mallen2022not and four multi-hop QA datasets, including HotpotQA yang2018hotpotqa , 2WikiMultiHopQA ho2020constructing , Bamboogle press2023measuring , and MuSiQue trivedi2022musique . Bamboogle and MuSiQue serve as out-of-domain QA evaluation datasets. The diversity of these datasets enables a comprehensive assessment of agentic RAG. We report Exact Match (EM) and F1 scores as evaluation metrics. Refer to Appendix E for more details about dataset introduction, statistics, and metrics.
Baselines. We implement 12 baseline models which can be categorized into 6 types as follows: Zero-shot: Directly use prompt engineering on LLM to answer the question without or with retrieved documents lewis2020retrieval . Active: Actively make additional retrieval when retrieved data or generated responses have low confidence jiang2023active ; asai2023self . Adaptive: Dynamically chooses the most suitable RAG pipeline from no-retrieval, single-hop, or multi-hop retrieval strategies jeong2024adaptive . RAG-CoT: Integrates chain-of-thought reasoning with retrieval, enabling multi-step, evidence-seeking answers trivedi2023interleaving ; shao2023enhancing . Summary: Compresses or summarizes retrieved content to fit model input constraints while retaining key information xu2024recomp ; jiang2024longllmlingua ; li2023compressing . Reasoning: Enhances multi-hop reasoning by structuring the reasoning process and scrutinizing retrieved evidence yu2024auto ; jin2025search . Note that ReasonRAG and all baselines use Qwen2.5-7B-Instruct yang2024qwen2 as the backbone model, ensuring fair comparison. Refer to more implementation details about ReasonRAG and baselines in Appendix F.1 and F.2.
@@ 3.3 Training Efficiency
Figure 4 compares the training efficiency of ReasonRAG and Search-R1, both trained on 4 NVIDIA A100 80GB GPUs. The figure illustrates the progression of EM scores with increasing GPU hours across three datasets. The results reveal that ReasonRAG has higher training efficiency compared to Search-R1. ReasonRAG achieves superior EM scores with fewer GPU hours, indicating that it requires less training data and compute to reach strong performance levels. In contrast, Search-R1 requires significantly more GPU hours to reach similar performance.
@@ 4 Related Works
Early prompt-based approaches leverage manually designed workflows to elicit the inherent capabilities of LLMs for interacting with external information. Specifically, the RAG task is often decomposed into subtasks such as adaptive retrieval judgment jeong2024adaptive , query generation amplayo2023query ; trivedi2023interleaving ; shao2023enhancing , evidence extraction asai2023self ; xu2024recomp ; li2023compressing ; jiang2024longllmlingua , and answer generation jiang2023active . Recently, agentic RAG aims to design workflows that empower LLM to autonomously interact with external information. OpenResearcher zheng2024openresearcher , AirRAG feng2025airrag , IterDRAG yue2024inference , PlanRAG verma2025plan , and search-o1 li2025search demonstrate strong performance improvement by the effective incorporation with the search engine. Nevertheless, these methods are limited by their dependence on inherent capabilities for interacting with external information and the requirement for manual design when applied to new domains.
@@
Outcome-supervised RL has rapidly emerged as a mainstream paradigm for enhancing LLM reasoning capabilities by providing end-to-end optimization. Following the widespread adoption of RL by major AI companies to improve the reasoning abilities of their models on complex tasks team2025kimi ; su2024dualformer ; shao2024deepseekmath , recent work jin2025search ; zheng2025deepresearcher ; wang2025otc ; sun2025zerosearch has extended outcome-supervised reinforcement learning to RAG, empowering LLMs to autonomously utilize search engines for intricate inference. While outcome-supervised RL has demonstrated significant performance gains, it also faces challenges such as reward sparsity, training instability, and substantial computational cost. Moreover, outcome-supervised RL typically demands extensive training resources. In contrast, process-supervised RL has recently been applied to enhance reasoning abilities, outperforming outcome-supervised approaches by providing fine-grained rewards ma2025s ; lin2025qlass ; wang2024math ; khalifa2025process . As an alternative avenue for improving LLM reasoning in RAG, process-supervised RL for RAG remains unexplored.
@@ 5 Conclusion
  • Appendix D: Visualizes our generated process-level reward annotations. This serves as an extended example for Section 2.2.3.
  • Appendix E: Provides additional details of the evaluation setup. This serves as a supplement to Section 3.1.
  • Appendix F: Provide more details of the implementation. This serves as a supplement to Section 3.1.
  • Appendix G: Details the prompt design for Agentic RAG. This serves a supplement to Section 2.4.
  • Appendix H: Details the licensing terms and conditions governing the use and distribution of the proposed datasets.
  • Appendix I: Discusses the limitations and constraints of the proposed approach.
  • Appendix J: Evaluates the potential societal implications and ethical considerations of the research.
@@ Appendix A Monte Carlo Tree Search for Process-level Reward
Based on the state, the policy of the next action is defined as
$\pi(a|s)=\text{LLM}(a|s)=\begin{cases}\pi_{\theta}(\cdot|x,y_{<i},p_{\text{stage}}),&\text{if stage is Reasoning}\\ \pi_{\theta}(\cdot|x,y_{<i},\text{docs},p_{\text{stage}}),&\text{otherwise}\\ \end{cases}$ (4)
@@
Selection: This step aims to select nodes that balance the search quality and exploration degree. The node selection starts from the root node, and iteratively selects the child nodes based on their state value $Q$ and visiting frequency $N$. These variable are refined during the search strategy, detailed in the backpropagation section. To effectively trade off between exploring unvisited nodes and exploiting nodes with higher state value, we iteratively search for the next node using UCT score [64]. The state is selected according to the following formula:

### v3

@@ Abstract
Retrieval-augmented generation (RAG) enhances the text generation capabilities of large language models (LLMs) by integrating external knowledge and up-to-date information. However, traditional RAG systems are limited by static workflows and lack the adaptability required for multistep reasoning and complex task management. To address these limitations, agentic RAG systems (e.g., DeepResearch) have been proposed, enabling dynamic retrieval strategies, iterative context refinement, and adaptive workflows for handling complex search queries beyond the capabilities of conventional RAG. Recent advances, such as Search-R1, have demonstrated promising gains using outcome-based reinforcement learning, where the correctness of the final answer serves as the reward signal. Nevertheless, such outcome-supervised agentic RAG methods face challenges including low exploration efficiency, gradient conflict, and sparse reward signals. To overcome these challenges, we propose to utilize fine-grained, process-level rewards to improve training stability, reduce computational costs, and enhance efficiency. Specifically, we introduce a novel method ReasonRAG that automatically constructs RAG-ProGuide, a high-quality dataset providing process-level rewards for (i) query generation, (ii) evidence extraction, and (iii) answer generation, thereby enhancing model inherent capabilities via process-supervised reinforcement learning. With the process-level policy optimization, the proposed framework empowers LLMs to autonomously invoke search, generate queries, extract relevant evidence, and produce final answers. Compared to existing approaches such as Search-R1 and traditional RAG systems, ReasonRAG, leveraging RAG-ProGuide, achieves superior performance on five benchmark datasets using only 5k training instances, significantly fewer than the 90k training instances required by Search-R1. Our code is available at https://github.com/Applied-Machine-Learning-Lab/ReasonRAG.
@@ 1 Introduction
Large language models (LLMs) brown2020language; chowdhery2023palm; achiam2023gpt have demonstrated substantial proficiency in text generation and natural language understanding zhularge; gonzalez2024does, revealing their potential for powering various downstream applications such as recommender systems zhang2025llmtreerec. However, their reliance on static training data constrains their ability to address dynamic and real-time queries, often resulting in outdated or hallucinated information ji2023survey; huang2025survey; mckenna2023sources. Retrieval-Augmented Generation (RAG) lewis2021RAG has emerged as a promising solution by equipping LLMs with external knowledge sources, improving the relevance, factual accuracy, and timeliness of responses gao2023retrieval; fan2024survey; xiong2024benchmarking. Despite these advances, traditional RAG architectures are limited by their linear and static workflows, which suffer from complex multi-step reasoning, deep contextual integration, and iterative response refinement lewis2021RAG; singh2025agentic. To address these shortcomings, agentic RAG (e.g., DeepResearch ds2024gemini; ds2025openai; zhang2025deep) systems have been developed, enabling dynamic retrieval strategies, enhanced contextual understanding, and iterative refinement. Achieving agentic RAG requires the underlying LLMs to orchestrate retrieval, filter relevant information, and iteratively refine their outputs, resulting in more adaptive and efficient information processing.
@@ 2.2.1 Shortest Path Reward Estimation (SPRE)
Formally, the agentic RAG process consists of an $n$-step sequence $y_{1},\cdots,y_{n}]$, where each $y_{i}$ represents the output of a single reasoning step, conditioned on the initial question $x$ and previous steps $y_{<i}$. To evaluate a partial sequence $y_{1:t}$, we simulate multiple continuations, known as rollouts, until a final answer is obtained. By repeating the rollout process $k$ times and scoring each outcome, we approximate the reward as a Monte Carlo-style estimation with step-based penalties:
@@ 2.2.2 Monte Carlo Tree Search (MCTS) for Process-level Exploration
2.2.3 RAG-ProGuide Dataset
Construction.
Using the MCTS-based exploration framework, we construct a high-quality process-supervised dataset RAG-ProGuide to facilitate process-level policy optimization. We randomly sample 3,000 questions each from PopQA mallen2022not, HotpotQA yang2018hotpotqa, and 2WikiMultihopQA ho2020constructing, covering both single-hop and multi-hop question answering tasks. An advanced large language model serves as the policy model to simulate the agentic RAG reasoning process within the MCTS framework (see Section 2.2.2). During tree search, we prune all branches that do not yield a final answer. For each complete trajectory, we compute the F1 score between the predicted answer and ground truth, and use this correctness signal to estimate intermediate node rewards via SPRE (see Section 2.2.1). These rewards are propagated through the MCTS tree to guide preference pair selection. To ensure high-quality preference data, we perform post-processing to remove duplicates and uninformative comparisons: (i) we discard preference pairs with identical response sequences, and (ii) pairs with a reward difference less than 0.01. After filtering, the final dataset consists of 4,603 questions and 13,289 distinct preference pairs. Figure 4 illustrates an example for the tree-structured process data. The root node corresponding to the original question, each node corresponds one-step response from LLMs. The correct reasoning path has been colored in orange and annotated with a higher reward.
Dataset Statistics.
Table 1 presents detailed statistics and distributions for our dataset. The questions are drawn from PopQA, HotpotQA, and 2WikiMultihopQA, providing comprehensive coverage of both single-hop and multi-hop reasoning within the RAG decision space. The dataset contains a balanced distribution of three reasoning actions, reflecting the multi-stage nature of the agentic RAG process. As shown in Figure 3(a), the distribution of preference pair types demonstrates diverse comparative scenarios; the x-axis abbreviations (A: answer generation, Q: query generation, E: evidence extraction) indicate action types in accepted versus rejected paths. This diversity ensures fine-grained comparative coverage across different reasoning stages. Figure 3(b) indicates a range of reasoning iteration counts, consistent with the complexity of multi-hop inference. Figure 3(c) shows a broad distribution of response token lengths, confirming the dataset’s capacity to capture various response complexities. Additionally, Figure 3(d) illustrates the probability distribution of the reward gap between preference pairs. The comprehensive coverage of this distribution across a wide range of values ensures that the preference learning is guided by a rich and informative signal. Collectively, these statistics demonstrate the dataset’s quality and its suitability for training robust process-level decision policies in agentic RAG frameworks.
@@ 2.3 Process-Supervised Preference Optimization
1:Initialize reasoning count $i\leftarrow 0$, and $\text{stage}\leftarrow\text{Reasoning}$
2:for $i\leftarrow 0$ to N1N-1 do
@@
19:end for
20:return final response $y$
@@ 2.4 Agentic RAG Inference
The workflow operates through three recursive decision states: Reasoning, Grounding, and Terminal. In the Reasoning state, the LLM evaluates the current context to decide if it has sufficient information to answer the question. If sufficient, it generates a final answer enclosed in placeholders (<answer> and </answer>), thus terminating the process. If not, the model creates a new query enclosed in <query> and </query> placeholders to retrieve additional evidence. The system then transitions to the Grounding state, where documents are retrieved based on the query, and the model extracts relevant evidence spans. These evidence spans are appended to the context, after which the process loops back to the Reasoning state for further deliberation. (See Appendix F for detailed prompt designs.)
@@ 3.1 Experimental Setup
Evaluation Dataset & Metrics. We evaluate ReasonRAG and all baselines on five public benchmarks: the single-hop QA dataset PopQA mallen2022not and four multi-hop QA datasets, including HotpotQA yang2018hotpotqa, 2WikiMultiHopQA ho2020constructing, Bamboogle press2023measuring, and MuSiQue trivedi2022musique. Bamboogle and MuSiQue serve as out-of-domain QA evaluation datasets. The diversity of these datasets enables a comprehensive assessment of agentic RAG. We report Exact Match (EM) and F1 scores as evaluation metrics. Refer to Appendix D for more details about dataset introduction, statistics, and metrics.
Baselines. We implement 12 baseline models which can be categorized into 6 types as follows: Zero-shot: Directly use prompt engineering on LLM to answer the question without or with retrieved documents lewis2020retrieval. Active: Actively make additional retrieval when retrieved data or generated responses have low confidence jiang2023active; asai2023self. Adaptive: Dynamically chooses the most suitable RAG pipeline from no-retrieval, single-hop, or multi-hop retrieval strategies jeong2024adaptive. RAG-CoT: Integrates chain-of-thought reasoning with retrieval, enabling multi-step, evidence-seeking answers trivedi2023interleaving; shao2023enhancing. Summary: Compresses or summarizes retrieved content to fit model input constraints while retaining key information xu2024recomp; jiang2024longllmlingua; li2023compressing. Reasoning: Enhances multi-hop reasoning by structuring the reasoning process and scrutinizing retrieved evidence yu2024auto; jin2025search; li2025search. Note that ReasonRAG and all baselines use Qwen2.5-7B-Instruct yang2024qwen2 as the backbone model, ensuring fair comparison. Refer to more implementation details about ReasonRAG and baselines in Appendix E.1 and E.2.
@@ 3.3 Training Efficiency
Figure 5 compares the training efficiency of ReasonRAG and Search-R1. The figure illustrates the progression of EM scores with increasing GPU hours across three datasets. The results reveal that ReasonRAG has higher training efficiency compared to Search-R1. ReasonRAG achieves superior EM scores with fewer GPU hours, indicating that it requires less training data and compute to reach strong performance levels. In contrast, Search-R1 requires significantly more GPU hours to reach similar performance.
@@ 4 Related Works
Early prompt-based approaches leverage manually designed workflows to elicit the inherent capabilities of LLMs for interacting with external information. Specifically, the RAG task is often decomposed into subtasks such as adaptive retrieval judgment jeong2024adaptive, query generation amplayo2023query; trivedi2023interleaving; shao2023enhancing, evidence extraction asai2023self; xu2024recomp; li2023compressing; jiang2024longllmlingua, and answer generation jiang2023active. While some efforts have focused on optimizing RAG through personalization li2025survey, graph-based retrieval xu2025align, or reranking techniques jia2025bridging, yet a critical gap remains in designing LLMs that can autonomously invoke search engines. Recently, agentic RAG aims to design workflows that empower LLM to autonomously interact with external information. OpenResearcher zheng2024openresearcher, AirRAG feng2025airrag, IterDRAG yue2024inference, PlanRAG verma2025plan, and Search-o1 li2025search demonstrate strong performance improvement by the effective incorporation with the search engine. Nevertheless, these methods are limited by their dependence on inherent capabilities for interacting with external information and the requirement for manual design when applied to new domains, and lack explicit mechanisms for eliminating distracting information 10806754.
@@
Reinforcement Learning (RL) has consistently delivered significant performance gains across a spectrum of sequential decision-making tasks, as evidenced by its successful application in domains such as recommender systems zhao2018deep; zhao2019deep; zhao2018recommendations; zhao2021dear; liu2023multi. Recently, the success of models like DeepSeek-R1 guo2025deepseek has vividly demonstrated the substantial potential of outcome-supervised RL in enhancing complex reasoning capabilities, establishing it as a mainstream paradigm for end-to-end optimization of LLMs. Following the widespread adoption of RL by major AI companies to improve the reasoning abilities of their models on complex tasks team2025kimi; su2024dualformer; shao2024deepseekmath, recent work jin2025search; zheng2025deepresearcher; wang2025otc; sun2025zerosearch has extended outcome-supervised reinforcement learning to RAG, empowering LLMs to autonomously utilize search engines for intricate inference. While outcome-supervised RL has demonstrated significant performance gains, it also faces challenges such as reward sparsity, training instability, and substantial computational cost. Moreover, outcome-supervised RL typically demands extensive training resources. In contrast, process-supervised RL has recently been applied to enhance reasoning abilities, outperforming outcome-supervised approaches by providing fine-grained rewards ma2025s; lin2025qlass; wang2024math; khalifa2025process. As an alternative avenue for improving LLM reasoning in RAG, process-supervised RL for RAG remains unexplored.
@@ 5 Conclusion
Acknowledgements
This research was partially supported by National Natural Science Foundation of China (No.62502404), Hong Kong Research Grants Council’s Research Impact Fund (No.R1015-23), Research Grants Council’s Collaborative Research Fund (No.C1043-24GF), Research Grants Council’s General Research Fund (No.11218325), Graduate Research Fund of the School of Economics and Management of Dalian University of Technology (No. DUTSEMDRFKO1), Institute of Digital Medicine of City University of Hong Kong (No.9229503), and Huawei (Huawei Innovation Research Program).
@@ Acknowledgements
  • Appendix D: Provides additional details of the evaluation setup. This serves as a supplement to Section 3.1.
  • Appendix E: Provide more details of the implementation. This serves as a supplement to Section 3.1.
  • Appendix F: Details the prompt design for Agentic RAG. This serves a supplement to Section 2.4.
  • Appendix G: Details the licensing terms and conditions governing the use and distribution of the proposed datasets.
  • Appendix H: Discusses the limitations and constraints of the proposed approach.
  • Appendix I: Evaluates the potential societal implications and ethical considerations of the research.
@@ Appendix A Monte Carlo Tree Search for Process-level Reward
Formally, for each RAG intermediate process, its corresponding state $s$ encompasses the original question $x$, the preceding thoughts $y_{<i}$, and a stage $\in\{\text{Reasoning, Grounding, Terminal}\}$. The stage indicates the current decision mode within the reasoning process. In the Reasoning stage, the LLM autonomously decides whether to generate a new query or directly produce an answer. If a query is chosen, it triggers a call to the external search engine, and the retrieved documents are added to the context in the next step. If the model opts to generate an answer instead, the process transitions into the Terminal stage. In the Grounding stage, the model extracts relevant evidence spans from the retrieved documents based on the most recent query. After extracting evidence, the state transitions back to the Reasoning stage, enabling further iterative reasoning.
Based on the state, the policy of the next action is defined as
@@
Selection: This step aims to select nodes that balance the search quality and exploration degree. The node selection starts from the root node, and iteratively selects the child nodes based on their state value $Q$ and visiting frequency $N$. These variable are refined during the search strategy, detailed in the backpropagation section. To effectively trade off between exploring unvisited nodes and exploiting nodes with higher state value, we iteratively search for the next node using UCT score [kocsis2006bandit]. The state is selected according to the following formula: